If x is a binomial random variable, compute p(x) for each of the following cases:

(a) n=5,x=5,p=0.2 p(x)=

(b) n=4,x=4,p=0.8 p(x)=
Answer #1

P(X = x) = nCx * px * (1 - p)n - x

a) P(X = 5) = 5C5 * (0.2)^5 * (0.8)^0 = 0.00032

b) P(X = 4) = 4C4 * (0.8)^4 * (0.2)^0 = 0.4096